### v1

# Improving Genetic Programming for Symbolic Regression with Equality Graphs

Abstract.
The search for symbolic regression models with genetic programming (GP) has a tendency of revisiting expressions in their original or equivalent forms. Repeatedly evaluating equivalent expressions is inefficient, as it does not immediately lead to better solutions. However, evolutionary algorithms require diversity and should allow the accumulation of inactive building blocks that can play an important role at a later point. The equality graph is a data structure capable of compactly storing expressions and their equivalent forms allowing an efficient verification of whether an expression has been visited in any of their stored equivalent forms. We exploit the e-graph to adapt the subtree operators to reduce the chances of revisiting expressions. Our adaptation, called eggp, stores every visited expression in the e-graph, allowing us to filter out from the available selection of subtrees all the combinations that would create already visited expressions. Results show that, for small expressions, this approach improves the performance of a simple GP algorithm to compete with PySR and Operon without increasing computational cost. As a highlight, eggp was capable of reliably delivering short and at the same time accurate models for a selected set of benchmarks from SRBench and a set of real-world datasets.
1. Introduction
Symbolic regression (SR) searches for a mathematical function that approximates a set of data points. It can be used to find nonlinear regression models (Koza, 1992; Kronberger et al., 2024a) or for scientific discovery (Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020; Russeil et al., 2024; Cao et al., [n. d.]; Lamb and Harrington, [n. d.]).
The current state-of-the-art SR algorithms (La Cava et al., 2021; de Franca et al., 2024) use genetic programming (GP) as the main search engine. One of such improvements, used by several state-of-the-art implementations, is the search for a parametric function that replaces the constants with adjustable parameters that are fitted using nonlinear optimization techniques.
Not only is the search for a symbolic model NP-hard (Virgolin and Pissis, [n. d.]) but when searching for a parametric model, it usually requires the solution to a multimodal optimization problem, which by itself is NP-hard (Murty and Kabadi, 1985) and can hinder the search for the optimal solution.
To make matters worse, the usual way of encoding mathematical expressions as symbolic expression trees, allows GP to visit different but equivalent expressions (Kronberger et al., 2024b) that evaluate to the same values. These equivalent expressions may be unnecessarily large and contain redundant parameters, which reduces the probability of finding their optimal values (de Franca and Kronberger, 2023; Kronberger and de França, 2024). Even for the simple expression we can produce an infinite number of equivalent expressions considering that are fitting parameters, for example , , , etc. are all different parameterizations of the same expression.
GP cannot differentiate between equivalent forms of a given expression, and simplification heuristics are often insufficient, as seen in (de Franca and Kronberger, 2023). Some authors (Keller and Banzhaf, 1999; Miller and Smith, 2006) argue that redundancy is necessary to allow the algorithm to navigate through the search space, as these equivalent expressions are guaranteed to have the same accuracy, allowing the search to keep multiple genetically different variations of solution candidates in the hopes of finding a better solutions. However, we do not know what would happen to GP search dynamics if we try to prevent keeping semantic duplicates in the GP population.
Equality saturation (Willsey et al., 2021) uses a set of equivalence rules and produces all equivalent forms of a given expression. It alleviates the phase ordering problem in the optimization phase when compiling computer programs. Given a program represented as a symbolic expression tree, and a set of equivalence rules, it iteratively applies the rules and stores all equivalent programs in a compact data structure called equality graph (e-graph). The main idea is that upon saturation, the graph contains all equivalent forms of the original program and the optimal form can be extracted from the e-graph using a heuristic. This technique was previously used in the context of SR in (de Franca and Kronberger, 2023; Kronberger and de França, 2024) to investigate the problem of overparameterization that can negatively affect the fitting of numerical parameters. The e-graph has another interesting feature that can be exploited by SR algorithms: it contains a database of all visited patterns and their equivalent forms that can be easily matched against new candidate expressions.
In this work, we investigate the benefits of exploiting the e-graph to increase the chances of generating a novel expression when applying crossover and mutation. In short, the proposed crossover operator will limit the selection of the second parent to a subset of subtrees that can only generate an unvisited expression. For the subtree mutation, it will replace a random subtree from the original tree with a random subtree generating the new tree . If has already been visited, the root of will be changed to a random choice of nonterminals that generates an unvisited expression.
These operators are tested inside a simple GP implementation based on tinyGP (Sipper, 2019) and compared against tinyGP with parameter fitting and two state-of-the-art algorithms: Operon (Burlacu et al., 2020b) and PySR (Cranmer, 2023). The results show that this simple111We acknowledge that the modification is simple provided we have a working implementation of e-graph. modification can improve the performance of tinyGP to an extent that it becomes competitive (and in some aspects better) than the state-of-the-art. The use of an e-graph as a support structure for GP brings new light to symbolic regression and GP with the possibility of exploring the accumulated history of the search process and even combining the history of multiple searches.
This paper is organized such that in Section 2 we will summarize the related works in symbolic regression. Section 3 will explain the basic concepts of equality saturation and the e-graph data structure. In Section 4 we will detail the proposed modification to simple GP. Section 5 will show the experiment methods used in this paper followed by Section 6 where we report and discuss the results. Finally, Section 7 gives some final remarks and expectations for the future.
2. Related work
The redundancy of GP search space has been investigated by many authors with conflicting conclusions to whether this is beneficial or not for the search. For example, Ebner (Ebner et al., 2001) argued that this redundancy enables the search to reach the optima through different trajectories, increasing the chances of achieving one of the equivalent expressions. On the other hand, Gustafson et al. (Gustafson et al., 2005) observed that when the recombination between two similar solutions was forbidden, there was an increase in offsprings that changed the original behavior of their parents, leading to increased performance.
Several works made a detailed study about the redundancy and neutrality in GP (i.e., when a change in the solution has no effect on its outcome). For example, Hu, Banzhaf, and Ochoa (Banzhaf et al., 2024; Hu and Banzhaf, 2018; Hu et al., 2023) investigated linear GP for Boolean SR problems with the help of search trajectory networks showing that some phenotypes are overrepresented in the search space. Regarding subtree crossover, McPhee et al. (McPhee et al., 2008) showed that over 75% of crossovers produced no immediately useful semantic changes.
Kronberger et al. (Kronberger et al., 2024b) studied the inefficiency of a simple GP comparing with the enumerated search space (Bartlett et al., 2023) and using equality saturation to count the percentage of unique expressions generated during the GP search. They found that from the total of visited expressions during the search, only around were unique. This not only wastes computational resource but it also shows that, at some point, GP fails to explore different regions of the search space.
Many authors observed improvements in the obtained solutions when applying any form of simplification during the search (Cao et al., 2023; Randall et al., 2022; Rivero et al., 2022; Seidyo Imai Aldeia et al., 2024) and, as a side effect, it stimulates the diversity of the population (Burlacu et al., 2019, 2020a).
Equality saturation has been used in the context of symbolic regression as a support tool to study the behavior of the search. Many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters (de Franca and Kronberger, 2023; Kronberger and de França, 2024). This redundancy can increase the chance of failing to correctly optimize such parameters, leading to sub-optimal solutions. In (Kronberger et al., 2024b) this technique was used to detect the equivalent expressions visited during the GP search. So far, the equality saturation technique was not used during the GP search to improve the quality of the solutions.
3. Equality saturation and e-graphs
Equality saturation (Tate et al., 2009) was proposed as a solution to the phase ordering problem in compiler optimization. This problem occurs when optimizing a program by applying a set of rewrite rules sequentially while dropping the information about the previous versions of the program. If the optimization follows a non-optimal sequence, it will lead to a sub-optimal program.
Equality saturation solves this issue by applying all of the optimization rules in parallel while keeping the intermediate transformations in a compact form using the data structure called e-graph.
Fig. 1a illustrates an example of an e-graph. Each solid box represents an e-node that contains a symbol of the expression. The dashed boxes group a set of e-nodes together and it is called an e-class, each e-class is assigned an e-class id (number in the bottom right of an e-class box). The main property of an e-class is that, no matter which e-node is chosen during the traversal, it will lead to an equivalent expression with all other e-nodes.
We can see that in the middle box (e-class id ), if we follow through it will generate the expression and if we follow through it will generate . The abstract description of the algorithm is very simple, though a concrete and optimal implementation requires the use of advanced techniques and data structures. The main idea is: i) match all the equivalence rules in the current state of the e-graph, ii) apply the rules creating new e-classes, iii) merge the equivalent e-classes, iv) repeat until saturation (i.e., no changes occur).
The current state-of-the-art implementation (Willsey et al., 2021)222https://docs.rs/egg/latest/egg/, called egg333egg stands for e-graphs good, encodes the e-graph as a map of an e-node to an e-class, a map of the e-class ids to their structure, and a set of e-classes that still needs analysis to maintain consistency. The structure of an e-class contains the information about the e-nodes it contains, a list of the parent e-nodes, and additional information also referred to as semantic analysis. This implementation also maintains a database of patterns that allows the algorithm to efficiently match patterns inside the e-graph structure.
Usually, the e-graph is used to represent a single program or expression and their equivalent forms. But, the structure can keep any number of expressions as long as we keep a list of the e-classes ids representing the root of each expression. For example, if we insert the expressions into the e-graph, we would end up with Fig. 1a, minus the equivalent relations. As a result, we would keep the list representing the ids of the expressions we inserted. New expressions are added bottom-up. Starting from a terminal, the algorithm checks whether it already exists in the e-graph and returns its e-class id, if it does not exist, it creates a new id. When adding an internal node, the algorithm first converts it to an e-node by replacing its children by their e-class ids and then it checks whether it already exists in the e-graph, returning the corresponding e-class id or a new one. In our example from Fig. 1, if we try to add the expression , it would first retrieve the e-class ids for the terminals and , then it would return the e-class id corresponding to the e-node , where underlined numbers correspond to e-class ids. Finally, it would create a new e-class with id and the e-node . This mechanism allows us to compactly store a set of expressions and readily assert whether an expression already exists in the structure.
In our implementation, we made a series of improvements to this data structure as we will describe in the next section.
4. eggp: e-graph GP
The proposed algorithm, eggp (e-graph genetic programming), follows the same structure as the traditional GP. The initial population is created using ramped half-and-half respecting a maximum size and maximum depth parameter (Koza, 1992) and, for a number of generations, it will choose two parents using tournament selection, apply the subtree crossover with probability followed by the subtree mutation with probability , when the offsprings replace the current population following a dominance criteria.
The key differences of eggp are:
new solutions are inserted into the e-graph followed by one step of equality saturation to find and store some of the equivalent expressions of the new offspring.
the current population is replaced by the set of individuals formed by: the Pareto front, the next front after excluding the first Pareto-front, and a selection of the last offspring at random until it reaches the desired population size.
the subtree crossover and mutation are modified to try to generate an unvisited expression.
Apart from that, we do not make use of other mutation operators commonly used in the literature (Burlacu et al., 2020b; Cranmer, 2023) or advanced techniques to stimulate diversity such as island models. Notice that a single step of equality saturation will not guarantee the insertion of all equivalent expressions in the e-graph but, as we apply more iterations, the e-graph can grow exponentially large. This issue is amplified by the fact that we are storing multiple expressions. As we will see in Sec. 6, the single step seems to be sufficient to improve the results and the benefits of increasing the number of steps is a subject for a future research. Regarding the Pareto-front extracted from the entire history, this is equivalent to the traditional NSGA-II algorithm (Deb et al., 2000) as the dominance relation is transitive. Keeping two ranks of dominance and filling up the remainder of the population with new expressions is meant to stimulate the combination of new expressions (exploration) while keeping the best fronts (exploitation). As a side-note, in this implementation we keep a database of generated expressions sorted by fitness and size (both objectives used in this work), so the Pareto-front can be retrieved in where is the number of extracted individuals. A new expression can be inserted into this structure in , where is the number of evaluated expressions so far. Finally, if we add the expression , equality saturation will generate the equivalent form (constants are replaced by parameters), this expression is stored in the database of expressions with size .
4.1. Crossover and mutation
Since the algorithm keeps the whole history of visited expression and its equivalent forms, we can exploit this information to increase the probability of generating unvisited expressions in the perturbation operators. For the crossover (Fig. 2a), the main idea is that it first chooses a crossover point from the first parent and then it randomly chooses a subtree of the second parent which generates an unvisited expression upon combination.
For the mutation operator (Fig. 2b), it chooses a point of the tree to generate a new subtree and, after generating the new subtree, it checks whether it generates an unvisited expression. With a negative answer, it randomly changes the root node of this subtree from a choice of all the possible non-terminals that can ensure novelty. In the event of impossibility of generating a new expression, a random recombination is returned.
To verify whether the combination of the original tree with the subtree replacing the -th node will generate an unvisited expression, we traverse the tree in pre-order traversal until we reach the node we want to replace (-th node). While traversing the tree, we build a list of -functions equivalent to a function that, given a tree , it creates a new tree with the operator of as the root, replacing one child with and keeping the other child of (if it is a binary operator). So, if the next step in the traversal is to the left, this function will build a tree rooted at with the left child being replaced by and the right child as , this is illustrated in Alg. 1. Upon reaching the desired node, we traverse the tree upwards by checking whether the current node exists in the e-graph and, if the assertion is false, it will apply the first function of the list to , generating the parent e-node, and calling this same function recursively (Alg. 2). Given that none of the ancestors exist in the e-graph, the algorithm confirms that this is an unvisited expression.
The full implementation of eggp has nine hyperparameters: number of generations, population size, maximum expression size, likelihood function (MSE, Gaussian, Poisson, Bernoulli, ROXY (Lelli et al., 2017)), number of iterations and retries for the parameter optimization, probabilities of crossover and mutation, and the list of non-terminals.
The algorithm is implemented in Haskell using the srtree444https://github.com/folivetti/srtree library for symbolic regression and with an adapted implementation of equality saturation based on hegg555https://github.com/alt-romes/hegg. The binaries are available at https://github.com/folivetti/srtree/releases/tag/v2.0.1.0.
5. Experiments
To measure the benefit of stimulating novelty using the history of visited expressions and their equivalent form stored in the e-graph, we chose three baseline algorithms: a version of tinyGP (Sipper, 2019) implemented using the same backend library, Operon (Burlacu et al., 2020b) and PySR (Cranmer, 2023).
Operon is a carefully crafted implementation of GP for symbolic regression with runtime performance in mind and a good set of default hyperparameters. It incorporates multiple mutation operators that allow a finer perturbation of a solution. Besides, it envelops every variable node with a scaling parameter adjusted using nonlinear optimization. PySR also supports the same mutation operators and nonlinear optimization of the parameters, it stands out with the use of an island model capable of keeping the diversity of the population to stimulate the exploration of the search space. It also applies a simplification heuristic on a selection of the expressions. Both PySR and Operon uses multi-objective optimization optimizing accuracy and size as default. The research questions we want to address with the experiments are:
What is the impact of increasing the probability of generating novelty, when compared to a minimalistic implementation such as tinyGP?
How close does eggp get to the state-of-the-art without resorting to more advanced concepts such as specialized mutation operators, enforcing the placement of numerical parameters, and island model?
We should stress that we have kept eggp with only the subtree crossover and mutation to be directly comparable with tinyGP, thus measuring the benefits of this modification.
We have fixed all the common hyperparameters to a standard value and performed a hyperparameter tuning with two different settings for each algorithm varying only whether to have a larger population with fewer iterations or a smaller population with more iterations. The only differences in settings are: for PySR we are using populations in its island model, so the size of each population is of the population size for the other algorithms, earlier experiments with the first set of benchmarks revealed that PySR performs significantly worse when using a single population; for Operon, we perform a maximum of optimization iterations instead of iterations with different starting points since it does not support multiple restarts for the parameter optimization; for eggp, of the training data is separated and used as a validation set to calculate the fitness, while the parameters are fitted using the remaining . Notice that tinyGP is the only single-objective approach and will be left out of the hypervolume comparison.
| Parameter | Value |
| number of evaluations | |
| prob. mutation | |
| prob. crossover | |
| non-terminal set | |
| pop. size / gens. / tourn. size | |
| max depth | |
| objectives | |
| optimization steps | ( for Operon) |
For the first set of experiments, we picked four datasets from the original SRBench (de Franca et al., 2024) and split each one into training and test sets with a ratio of . We executed each configuration times and stored the test errors of each run. Once we have the optimal hyperparameters, we evaluate these algorithms in the benchmarks of the current version of SRBench666https://github.com/cavalab/srbench/discussions/174#discussioncomment-10285133 for the reduced SRBench track. This set is supposed to be representative as it contains datasets with different characteristics. For this experiment, we applied a -fold cross-validation repeating the experiment times, creating a total of runs. Finally, we picked some real-world datasets from the literature corresponding to data from different fields, these data already have pre-determined training and test sets, as such we will run repetitions of each experiment. For every experiment, we store the final Pareto front and will report the performance plot, the AUC and average rank among all datasets and statistical test of the ranks, the average and standard deviation of the running time, the hypervolume of the Pareto front. All algorithms were restricted to a single core to ensure equal conditions.
Table 2 shows the datasets of each set of experiment with their corresponding number of data points, features and the chosen maximum size parameter. 777The maximum size for Operon is set to of the maximum size because, internally, Operon will not count the scale coefficients of a terminal towards the model size. This factor enforces Operon to search on a similar search space as the other algorithms.
| Name | Points | Features | max. size |
| hyperparameter tuning | |||
| 523_analcatdata_neavote | |||
| 527_analcatdata_election2000 | |||
| 635_fri_c0_250_10 | |||
| 1029_LEV | |||
| SRBench | |||
| 192_vineyard | |||
| 210_cloud | |||
| 522_pm10 | |||
| 557_analcatdata_apnea1 | |||
| 579_fri_c0_250_5 | |||
| 606_fri_c2_1000_10 | |||
| 650_fri_c0_500_50 | |||
| 678_visualizing_environmental | |||
| 1028_SWD | |||
| 1089_USCrime | |||
| 1193_BNG_lowbwt | |||
| 1199_BNG_echoMonths | |||
| Real world | |||
| Chemical_1_tower | |||
| Chemical_2_competition | |||
| Friction_stat_one-hot | |||
| Friction_dyn_one-hot | |||
| Flow_stress_phip0.1 | |||
| Nasa_battery_1_10min | |||
| Nasa_battery_2_20min | |||
| Nikuradse_1 | |||
| Nikuradse_2 | |||
6. Results and Discussion
6.1. Hyperparameter tuning
Applying the Wilcoxon rank test to the obtained results using , we observed that only tinyGP expressed a significant difference in their results (-value ) favoring generations and population size . There was no significant difference observed in any algorithm, thus, we chose the configuration which found the best result for most datasets. With this criteria, PySR and Operon was set to generations and eggp with generations.
6.2. New SRBench datasets
In Fig. 3 we can see the performance plots for each SRBench dataset of the best solution according to the highest on the training set. The x-axis of these plots represents the measured on the test set, and the y-axis shows the percentage of runs that the algorithm found an equal or larger than . The ideal algorithm would cover the whole area from to . As we can see from these plots, eggp always covers an area similar or better than the competing algorithms. Considering the area under the curve values (AUC), eggp covered the most area in out of the datasets, Operon in , PySR in and tinyGP in . It is also evident from these plots how eggp displays a significant improvement over tinyGP.
In Table 3 we can see the ranks when considering the median of the test set and the AUC. We can see that, using this criteria, eggp is consistently ranked second, while Operon and PySR take turns in the first and third place. The statistical test reveals that we can reject the null hypothesis when comparing to PySR with the alternative of having greater median rank. On the other hand, for the AUC values, we can see that eggp is greater than the other algorithms on average while rejecting the null hypotheses for each comparison. Unlike the median of the , the AUC is the average weighted by the probability of obtaining that value or greater, acting as a reliability measure of obtaining that value or greater.
| dataset | eggp | Operon | PySR | tinyGP |
| 192_vineyard | 1 | 4 | 3 | 2 |
| 210_cloud | 1 | 4 | 2 | 3 |
| 522_pm10 | 3 | 4 | 1 | 2 |
| 557_analcatdata_apnea1 | 2 | 4 | 3 | 1 |
| 579_fri_c0_250_5 | 2 | 1 | 3 | 4 |
| 606_fri_c2_1000_10 | 1 | 2 | 3 | 4 |
| 650_fri_c0_500_50 | 2 | 1 | 3 | 4 |
| 678_visualizing_environmental | 1 | 4 | 2 | 3 |
| 1028_SWD | 4 | 1 | 2 | 3 |
| 1089_USCrime | 1 | 3 | 4 | 2 |
| 1193_BNG_lowbwt | 2 | 1 | 3 | 4 |
| 1199_BNG_echoMonths | 4 | 3 | 1 | 2 |
| mean | 2.00 | 2.67 | 2.50 | 2.83 |
| -value | ||||
| 192_vineyard | 0.24 | 0.08 | 0.16 | 0.17 |
| 210_cloud | 0.71 | 0.34 | 0.71 | 0.58 |
| 522_pm10 | 0.16 | 0.14 | 0.19 | 0.18 |
| 557_analcatdata_apnea1 | 0.75 | 0.40 | 0.58 | 0.68 |
| 579_fri_c0_250_5 | 0.91 | 0.95 | 0.83 | 0.79 |
| 606_fri_c2_1000_10 | 0.96 | 0.96 | 0.82 | 0.70 |
| 650_fri_c0_500_50 | 0.87 | 0.92 | 0.83 | 0.40 |
| 678_visualizing_environmental | 0.28 | 0.08 | 0.27 | 0.20 |
| 1028_SWD | 0.38 | 0.39 | 0.39 | 0.36 |
| 1089_USCrime | 0.68 | 0.59 | 0.59 | 0.59 |
| 1193_BNG_lowbwt | 0.56 | 0.57 | 0.56 | 0.55 |
| 1199_BNG_echoMonths | 0.40 | 0.30 | 0.43 | 0.35 |
| mean | 0.58 | 0.48 | 0.53 | 0.46 |
| -value |
6.3. Real-world datasets
In Fig. 4 we observe a similar behavior with eggp covering an area close to or better than the best competing algorithm. Considering AUC, it found out of best results, while PySR only , Operon and tinyGP . Again, considering the ranks on the median of the (Table 4), it consistently achieved second place, but in this set, Operon was ranked first for most of the datasets. We should notice, though, that eggp results were always close to the best competing algorithm, while even Operon misbehaved in two datasets (flow and niku-2). When observing the statistical test results, we can conclude that there are no significant differences between the eggp and Operon, but the hypotheses of eggp being equivalent to PySR and tinyGP can be rejected.
| dataset | eggp | Operon | PySR | tinyGP |
| Chemical_1_tower | 2 | 1 | 4 | 3 |
| Chemical_2_competition | 2 | 1 | 4 | 3 |
| Flow_stress_phip0.1 | 1 | 4 | 2 | 3 |
| Friction_dyn_one-hot | 2 | 1 | 4 | 3 |
| Friction_stat_one-hot | 2 | 1 | 4 | 3 |
| Nasa_battery_1_10min | 2 | 1 | 4 | 3 |
| Nasa_battery_2_20min | 2 | 3 | 4 | 1 |
| Nikuradse_1 | 2 | 1 | 4 | 3 |
| Nikuradse_2 | 3 | 1 | 4 | 2 |
| mean | 2.00 | 1.56 | 3.78 | 2.67 |
| -value | ||||
| Chemical_1_tower | 0.86 | 0.89 | 0.67 | 0.81 |
| Chemical_2_competition | 0.58 | 0.66 | 0.07 | 0.37 |
| Flow_stress_phip0.1 | 0.99 | 0.51 | 0.98 | 0.88 |
| Friction_dyn_one-hot | 0.81 | 0.85 | 0.66 | 0.74 |
| Friction_stat_one-hot | 0.77 | 0.82 | 0.30 | 0.65 |
| Nasa_battery_1_10min | 0.96 | 0.96 | 0.85 | 0.93 |
| Nasa_battery_2_20min | 0.97 | 0.95 | 0.97 | 0.98 |
| Nikuradse_1 | 0.85 | 0.89 | 0.62 | 0.75 |
| Nikuradse_2 | 0.98 | 0.83 | 0.95 | 0.95 |
| mean | 0.87 | 0.82 | 0.67 | 0.79 |
| -value |
6.4. Hypervolume
Table 5 shows the average and standard deviation of the hypervolume of the final Pareto fronts. For SRBench, eggp covered a similar area to the other algorithms. PySR returned the largest hypervolume in most of the SRBench datasets. For the real-world datasets, eggp covered the most hypervolume overall.
| dataset | egraphGP | operon | PySR |
| srbench | |||
| 192 | |||
| 210 | |||
| 522 | |||
| 557 | |||
| 579 | |||
| 606 | |||
| 650 | |||
| 678 | |||
| 1028 | |||
| 1089 | |||
| 1193 | |||
| 1199 | |||
| real-world | |||
| Chemical_1 | |||
| Chemical_2 | |||
| Flow_stress | |||
| Friction_dyn | |||
| Friction_stat | |||
| Nasa_1 | |||
| Nasa_2 | |||
| Nikuradse_1 | |||
| Nikuradse_2 | |||
6.5. Computational time
Regarding the runtime, since Operon is the fastest symbolic regression implementation, as noted in (Burlacu et al., 2020b), we calculated the ratio between the average runtime of each algorithm to Operon. Fig. 5 shows the relative runtime per dataset. In this plot we can see that eggp and tinyGP were both between to times slower than Operon. PySR varied from to times the runtime of Operon depending on the dataset. The higher ratios were observed on high-dimensional or larger datasets. We should stress that all algorithms were constrained to run with a single thread, thus both Operon and PySR runtime could be smaller when exploiting multi-threading.
7. Conclusions
In this paper we explored the use of e-graphs and equality saturation as a mechanism to keep track of the history of the search engine of symbolic regression and exploiting its pattern matching capabilities to propose a variation to the traditional subtree crossover and mutation that increases the probability of generating novel expressions.
The e-graph data structure efficiently stores and queries for parts of expressions. Exploiting this capability, we modified the subtree operators to only sample subtrees that would form unvisited expressions. The expectation is that this simple modification would render a significant improvement in the search procedure.
We tested the proposed algorithm, called eggp, in different benchmarks from the literature and compared with a simple GP with the original subtree operators, and two state-of-the-art algorithms, PySR and Operon. The results showed that the modified operators are capable of improving the performance of a simple GP to compete with the state-of-the-art. The main highlight of this approach is that it consistently performs equal or better than the best competing approach.
The main limitations of these experiments lie in the use of default or reasonable values for the hyperparameters. We should notice that eggp contains eight parameters that should be adjusted accordingly in a practical scenario. Meanwhile, PySR contains about hyperparameters that may affect the algorithm performance888we are not considering hyperparameters that expect prior knowledge. and Operon contains about hyperparameters. With careful experimentation, both PySR and Operon could possibly achieve similar results. Regarding the runtime, eggp is consistently faster than PySR but significantly slower than Operon. When comparing with tinyGP, we can see that the use of e-graph and equality saturation does not increase the runtime significantly.
In conclusion, the expressiveness and capabilities of the e-graph data structure enabled us to make a simple modification to the original subtree operators while significantly improving the performance of GP for symbolic regression. This allowed us to obtain a more robust algorithm delivering better performance more reliably than state-of-the-art implementations. As for the next steps, this same modifications can be applied to any other mutation operator used by the state-of-the-art algorithms. Not only that, but the e-graph opens up many new possibilities for improving the search as it allows us to query expressions with a combination of properties, which can translate to diversity-preserving population and easy to integrate prior knowledge (Kronberger et al., 2022; Reuter et al., 2024). In addition, the storage of the search history allows us to analyze the learned building blocks and exploit this information to generate new solutions. 999The authors proudly made no use of LLMs for this work.
Acknowledgements.
F.O.F. is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) grant 301596/2022-0. G.K. is supported by the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology, the Federal Ministry for Labour and Economy, and the regional government of Upper Austria within the COMET project ProMetHeus (904919) supported by the Austrian Research Promotion Agency (FFG). No LLMs were used for this work.References
- (1)
- Banzhaf et al. (2024) Wolfgang Banzhaf, Ting Hu, and Gabriela Ochoa. 2024. How the Combinatorics of Neutral Spaces Leads Genetic Programming to Discover Simple Solutions. In Genetic Programming Theory and Practice XX. Springer, 65–86.
- Bartlett et al. (2023) Deaglan J Bartlett, Harry Desmond, and Pedro G Ferreira. 2023. Exhaustive symbolic regression. IEEE Transactions on Evolutionary Computation (2023).
- Burlacu et al. (2019) Bogdan Burlacu, Michael Affenzeller, Gabriel Kronberger, and Michael Kommenda. 2019. Online diversity control in symbolic regression via a fast hash-based tree similarity measure. In 2019 IEEE congress on evolutionary computation (CEC). IEEE, 2175–2182.
- Burlacu et al. (2020a) Bogdan Burlacu, Lukas Kammerer, Michael Affenzeller, and Gabriel Kronberger. 2020a. Hash-based tree similarity and simplification in genetic programming for symbolic regression. In Computer Aided Systems Theory–EUROCAST 2019: 17th International Conference, Las Palmas de Gran Canaria, Spain, February 17–22, 2019, Revised Selected Papers, Part I 17. Springer, 361–369.
- Burlacu et al. (2020b) Bogdan Burlacu, Gabriel Kronberger, and Michael Kommenda. 2020b. Operon C++ an efficient genetic programming framework for symbolic regression. In Proceedings of the 2020 Genetic and Evolutionary Computation Conference Companion. 1562–1570.
- Cao et al. ([n. d.]) Haotian Cao, Garrett W Merz, Kyle Cranmer, and Gary Shiu. [n. d.]. Learning Conformal Field Theory with Symbolic Regression: Recovering the Symbolic Expressions for the Energy Spectrum. ([n. d.]).
- Cao et al. (2023) Lulu Cao, Zimo Zheng, Chenwen Ding, Jinkai Cai, and Min Jiang. 2023. Genetic Programming Symbolic Regression with Simplification-Pruning Operator for Solving Differential Equations. In International Conference on Neural Information Processing. Springer, 287–298.
- Cranmer (2023) Miles Cranmer. 2023. Interpretable Machine Learning for Science with PySR and SymbolicRegression.jl. https://doi.org/10.48550/ARXIV.2305.01582
- de Franca and Kronberger (2023) Fabricio Olivetti de Franca and Gabriel Kronberger. 2023. Reducing Overparameterization of Symbolic Regression Models with Equality Saturation. In Proceedings of the Genetic and Evolutionary Computation Conference. 1064–1072.
- de Franca et al. (2024) Fabricio O de Franca, Marco Virgolin, M Kommenda, MS Majumder, M Cranmer, G Espada, L Ingelse, A Fonseca, M Landajuela, B Petersen, et al. 2024. SRBench++: Principled benchmarking of symbolic regression with domain-expert interpretation. IEEE transactions on evolutionary computation (2024).
- Deb et al. (2000) Kalyanmoy Deb, Samir Agrawal, Amrit Pratap, and Tanaka Meyarivan. 2000. A fast elitist non-dominated sorting genetic algorithm for multi-objective optimization: NSGA-II. In Parallel Problem Solving from Nature PPSN VI: 6th International Conference Paris, France, September 18–20, 2000 Proceedings 6. Springer, 849–858.
- Ebner et al. (2001) Marc Ebner, Mark Shackleton, and Rob Shipman. 2001. How neutral networks influence evolvability. Complex. 7, 2 (Nov. 2001), 19–33. https://doi.org/10.1002/cplx.10021
- Gustafson et al. (2005) Steven Gustafson, Edmund K Burke, and Natalio Krasnogor. 2005. On improving genetic programming for symbolic regression. In 2005 IEEE Congress on Evolutionary Computation, Vol. 1. IEEE, 912–919.
- Hu and Banzhaf (2018) Ting Hu and Wolfgang Banzhaf. 2018. Neutrality, robustness, and evolvability in genetic programming. Genetic Programming Theory and Practice XIV (2018), 101–117.
- Hu et al. (2023) Ting Hu, Gabriela Ochoa, and Wolfgang Banzhaf. 2023. Phenotype search trajectory networks for linear genetic programming. In European Conference on Genetic Programming (Part of EvoStar). Springer, 52–67.
- Keller and Banzhaf (1999) Robert E Keller and Wolfgang Banzhaf. 1999. The evolution of genetic code in genetic programming. In Proceedings of the Genetic and Evolutionary Computation Conference, Vol. 2. Morgan Kaufmann Orlando, 1077–1082.
- Koza (1992) John R. Koza. 1992. Genetic Programming: On the Programming of Computers by Means of Natural Selection. MIT Press, Cambridge, MA, USA.
- Kronberger et al. (2024a) Gabriel Kronberger, Bogdan Burlacu, Michael Kommenda, Stephan M. Winkler, and Michael Affenzeller. 2024a. Symbolic Regression. Chapman & Hall / CRC Press.
- Kronberger and de França (2024) Gabriel Kronberger and Fabrício Olivetti de França. 2024. Effects of reducing redundant parameters in parameter optimization for symbolic regression using genetic programming. Journal of Symbolic Computation (2024), 102413.
- Kronberger et al. (2022) Gabriel Kronberger, Fabricio Olivetti de França, Bogdan Burlacu, Christian Haider, and Michael Kommenda. 2022. Shape-constrained symbolic regression—improving extrapolation with prior knowledge. Evolutionary Computation 30, 1 (2022), 75–98.
- Kronberger et al. (2024b) Gabriel Kronberger, Fabricio Olivetti de Franca, Harry Desmond, Deaglan J. Bartlett, and Lukas Kammerer. 2024b. The Inefficiency of Genetic Programming for Symbolic Regression. In Parallel Problem Solving from Nature – PPSN XVIII, Michael Affenzeller, Stephan M. Winkler, Anna V. Kononova, Heike Trautmann, Tea Tušar, Penousal Machado, and Thomas Bäck (Eds.). Springer Nature Switzerland, Cham, 273–289.
- La Cava et al. (2021) William La Cava, Bogdan Burlacu, Marco Virgolin, Michael Kommenda, Patryk Orzechowski, Fabrício Olivetti de França, Ying Jin, and Jason H Moore. 2021. Contemporary symbolic regression methods and their relative performance. Advances in neural information processing systems 2021, DB1 (2021), 1.
- Lamb and Harrington ([n. d.]) Kara D Lamb and Jerry Y Harrington. [n. d.]. Discovering How Ice Crystals Grow Using Neural ODE’s and Symbolic Regression. ([n. d.]).
- Lelli et al. (2017) Federico Lelli, Stacy S. McGaugh, James M. Schombert, and Marcel S. Pawlowski. 2017. One Law to Rule Them All: The Radial Acceleration Relation of Galaxies. Astrophysical Journal 836, 2, Article 152 (Feb. 2017), 152 pages. https://doi.org/10.3847/1538-4357/836/2/152 arXiv:1610.08981 [astro-ph.GA]
- McPhee et al. (2008) Nicholas Freitag McPhee, Brian Ohs, and Tyler Hutchison. 2008. Semantic building blocks in genetic programming. In Genetic Programming: 11th European Conference, EuroGP 2008, Naples, Italy, March 26-28, 2008. Proceedings 11. Springer, 134–145.
- Miller and Smith (2006) J.F. Miller and S.L. Smith. 2006. Redundancy and computational efficiency in Cartesian genetic programming. IEEE Transactions on Evolutionary Computation 10, 2 (2006), 167–174. https://doi.org/10.1109/TEVC.2006.871253
- Murty and Kabadi (1985) Katta G Murty and Santosh N Kabadi. 1985. Some NP-complete problems in quadratic and nonlinear programming. Technical Report.
- Randall et al. (2022) David L Randall, Tyler S Townsend, Jacob D Hochhalter, and Geoffrey F Bomarito. 2022. Bingo: a customizable framework for symbolic regression with genetic programming. In Proceedings of the Genetic and Evolutionary Computation Conference Companion. 2282–2288.
- Reuter et al. (2024) Julia Reuter, Viktor Martinek, Roland Herzog, and Sanaz Mostaghim. 2024. Unit-Aware Genetic Programming for the Development of Empirical Equations. In International Conference on Parallel Problem Solving from Nature. Springer, 168–183.
- Rivero et al. (2022) Daniel Rivero, Enrique Fernandez-Blanco, and Alejandro Pazos. 2022. DoME: A deterministic technique for equation development and Symbolic Regression. Expert Systems with Applications 198 (2022), 116712.
- Russeil et al. (2024) Etienne Russeil, Fabrício Olivetti de França, Konstantin Malanchev, Bogdan Burlacu, Emille Ishida, Marion Leroux, Clément Michelin, Guillaume Moinard, and Emmanuel Gangler. 2024. Multiview Symbolic Regression. In Proceedings of the Genetic and Evolutionary Computation Conference. 961–970.
- Seidyo Imai Aldeia et al. (2024) Guilherme Seidyo Imai Aldeia, Fabricio Olivetti de Franca, and William G La Cava. 2024. Inexact Simplification of Symbolic Regression Expressions with Locality-sensitive Hashing. arXiv e-prints (2024), arXiv–2404.
- Sipper (2019) M. Sipper. 2019. Tiny Genetic Programming in Python. https://github.com/moshesipper/tiny_gp.
- Tate et al. (2009) Ross Tate, Michael Stepp, Zachary Tatlock, and Sorin Lerner. 2009. Equality saturation: a new approach to optimization. In Proceedings of the 36th annual ACM SIGPLAN-SIGACT symposium on Principles of programming languages. 264–276.
- Udrescu and Tegmark (2020) Silviu-Marian Udrescu and Max Tegmark. 2020. AI Feynman: A physics-inspired method for symbolic regression. Science Advances 6, 16 (2020), eaay2631.
- Virgolin and Pissis ([n. d.]) Marco Virgolin and Solon P Pissis. [n. d.]. Symbolic Regression is NP-hard. Transactions on Machine Learning Research ([n. d.]).
- Willsey et al. (2021) Max Willsey, Chandrakana Nandi, Yisu Remy Wang, Oliver Flatt, Zachary Tatlock, and Pavel Panchekha. 2021. Egg: Fast and extensible equality saturation. Proceedings of the ACM on Programming Languages 5, POPL (2021), 1–29.

### v2

@@
 1. Introduction
-Symbolic regression (SR) searches for a mathematical function that approximates a set of data points.
-It can be used to find nonlinear regression models (Koza, 1992; Kronberger et al., 2024a) or for scientific discovery (Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020; Russeil et al., 2024; Cao et al., [n.
-d.]; Lamb and Harrington, [n.
-d.]).
-The current state-of-the-art SR algorithms (La Cava et al., 2021; de Franca et al., 2024) use genetic programming (GP) as the main search engine.
-One of such improvements, used by several state-of-the-art implementations, is the search for a parametric function that replaces the constants with adjustable parameters that are fitted using nonlinear optimization techniques.
-Not only is the search for a symbolic model NP-hard (Virgolin and Pissis, [n.
-d.]) but when searching for a parametric model, it usually requires the solution to a multimodal optimization problem, which by itself is NP-hard (Murty and Kabadi, 1985) and can hinder the search for the optimal solution.
-To make matters worse, the usual way of encoding mathematical expressions as symbolic expression trees, allows GP to visit different but equivalent expressions (Kronberger et al., 2024b) that evaluate to the same values.
-These equivalent expressions may be unnecessarily large and contain redundant parameters, which reduces the probability of finding their optimal values (de Franca and Kronberger, 2023; Kronberger and de França, 2024).
-Even for the simple expression we can produce an infinite number of equivalent expressions considering that are fitting parameters, for example , , , etc.
-are all different parameterizations of the same expression.
-GP cannot differentiate between equivalent forms of a given expression, and simplification heuristics are often insufficient, as seen in (de Franca and Kronberger, 2023).
-Some authors (Keller and Banzhaf, 1999; Miller and Smith, 2006) argue that redundancy is necessary to allow the algorithm to navigate through the search space, as these equivalent expressions are guaranteed to have the same accuracy, allowing the search to keep multiple genetically different variations of solution candidates in the hopes of finding a better solutions.
-However, we do not know what would happen to GP search dynamics if we try to prevent keeping semantic duplicates in the GP population.
-Equality saturation (Willsey et al., 2021) uses a set of equivalence rules and produces all equivalent forms of a given expression.
-It alleviates the phase ordering problem in the optimization phase when compiling computer programs.
-Given a program represented as a symbolic expression tree, and a set of equivalence rules, it iteratively applies the rules and stores all equivalent programs in a compact data structure called equality graph (e-graph).
-The main idea is that upon saturation, the graph contains all equivalent forms of the original program and the optimal form can be extracted from the e-graph using a heuristic.
-This technique was previously used in the context of SR in (de Franca and Kronberger, 2023; Kronberger and de França, 2024) to investigate the problem of overparameterization that can negatively affect the fitting of numerical parameters.
-The e-graph has another interesting feature that can be exploited by SR algorithms: it contains a database of all visited patterns and their equivalent forms that can be easily matched against new candidate expressions.
-In this work, we investigate the benefits of exploiting the e-graph to increase the chances of generating a novel expression when applying crossover and mutation.
-In short, the proposed crossover operator will limit the selection of the second parent to a subset of subtrees that can only generate an unvisited expression.
-For the subtree mutation, it will replace a random subtree from the original tree with a random subtree generating the new tree .
-If has already been visited, the root of will be changed to a random choice of nonterminals that generates an unvisited expression.
-These operators are tested inside a simple GP implementation based on tinyGP (Sipper, 2019) and compared against tinyGP with parameter fitting and two state-of-the-art algorithms: Operon (Burlacu et al., 2020b) and PySR (Cranmer, 2023).
-The results show that this simple111We acknowledge that the modification is simple provided we have a working implementation of e-graph.
-modification can improve the performance of tinyGP to an extent that it becomes competitive (and in some aspects better) than the state-of-the-art.
+Symbolic regression (SR) (Koza, 1992; Kronberger et al., 2024a) searches for a mathematical function that approximates a set of data points often used for scientific discovery (Schmidt and Lipson, 2009; Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020; Russeil et al., 2024; Cao et al., 2024).
+The current SotA (La Cava et al., 2021; de Franca et al., 2024) uses genetic programming (GP) as the main search engine incorporating numerical parameters that can be fitted into the data using optimization techniques.
+The search for a symbolic model is NP-hard (Virgolin and Pissis, 2022) and when searching for a parametric model, it also requires the solution to a multimodal optimization problem, which by itself is NP-hard (Murty and Kabadi, 1985) and can hinder the search for the optimal solution.
+To make matters worse, the usual way of encoding mathematical expressions as symbolic expression trees, allows GP to visit semantically equivalent expressions111in this paper, we will refer to semantically equivalence as simply equivalence.
+with different syntax (Kronberger et al., 2024b).
+These equivalent expressions may be unnecessarily large and with redundant parameters, reducing the probability of finding their optimal values (de Franca and Kronberger, 2023; Kronberger and Olivetti de França, 2025).
+Even for the simple expression we can produce an infinite number of equivalent expressions considering that are fitting parameters, for example , , , are all different parameterizations of the same expression.
+GP cannot easily differentiate between equivalent expressions, and applying simplification heuristics are often insufficient, as seen in (de Franca and Kronberger, 2023).
+Some authors (Keller and Banzhaf, 1999; Miller and Smith, 2006) argue that redundancy is necessary to allow the algorithm to navigate through the search space, as these equivalent expressions are guaranteed to have the same accuracy, allowing the search to keep multiple genetically different variations of solution candidates in the hopes of finding better solutions.
+However, this supposition could not be validate as there were no means to efficiently verify equivalence.
+Equality saturation (Willsey et al., 2021) can produce all the equivalent expressions of a given expression through the parallel application of a set of equivalence rules.
+Given an expression represented as a directed acyclic graph, and a set of equivalence rules, it iteratively applies the rules and stores all equivalent programs in a compact data structure called equality graph (e-graph).
+The main idea is that upon saturation, the graph will contain all equivalent forms of the original program and the optimal form can be extracted from the e-graph using a heuristic cost function.
+This technique was previously used in the context of SR in (de Franca and Kronberger, 2023; Kronberger and Olivetti de França, 2025) to investigate the problem of overparameterization that can negatively affect the fitting of numerical parameters.
+The e-graph has another feature that can be exploited by SR algorithms: it implements an efficient pattern matching algorithm that can answer whether a given expression, or any of its equivalents, are already stored in the e-graph structure.
+In this work we introduce eggp, a GP algorithm that exploits the pattern matching capabilities of the e-graph to try to enforce the generation of unvisited expressions when applying the crossover and mutation operators.
+In this context, unvisited expressions mean any expression, or their equivalents, that was not previously evaluated during the history of search.
+In short, after choosing the crossover point of the first parent, the choices of points of the second parent are limited to those that will ensure the generation of an unvisited expression.
+For the mutation, after choosing a node of the expression at random, it will generate a new subtree, limiting the choice of its root node to the set that will ensure the generation of an unvisited expression.
+This procedure will enforce the introduction of novel solutions not only w.r.t.
+the current population but to the entire history of the search.
+The research questions we want to address in this paper are:
+What is the impact of increasing the probability of generating novelty, when compared to a minimalist implementation of GP for SR?
+How close does eggp get to the state-of-the-art without resorting to more advanced concepts such as specialized mutation operators, enforcing the placement of numerical parameters, and promoting diversity throug island model?
+These operators are tested inside a minimalist GP implementation, and compared against this same algorithm with the original subtree operators, and two high performant algorithms: Operon (Burlacu et al., 2020b) and PySR (Cranmer, 2023).
+The results show that this simple modification, provided we have a working implementation of the e-graph, can improve the performance of this minimalist GP to an extent that it becomes competitive (and in some aspects better) than the state-of-the-art.
 The use of an e-graph as a support structure for GP brings new light to symbolic regression and GP with the possibility of exploring the accumulated history of the search process and even combining the history of multiple searches.
@@
 Section 3 will explain the basic concepts of equality saturation and the e-graph data structure.
-In Section 4 we will detail the proposed modification to simple GP.
-Section 5 will show the experiment methods used in this paper followed by Section 6 where we report and discuss the results.
-Finally, Section 7 gives some final remarks and expectations for the future.
+In Section 4 we will detail the proposed modifications to the subtree operators.
+Section 5 and 6 will respectively detail the experiment methods, report and discuss the results.
+Finally, Section 7 will provide some final remarks and expectations for the future.
 2. Related work
@@
 Kronberger et al.
-(Kronberger et al., 2024b) studied the inefficiency of a simple GP comparing with the enumerated search space (Bartlett et al., 2023) and using equality saturation to count the percentage of unique expressions generated during the GP search.
+(Kronberger et al., 2024b) studied the inefficiency of a simple GP comparing with the enumerated search space (Bartlett et al., 2024) and using equality saturation to count the percentage of unique expressions generated during the GP search.
 They found that from the total of visited expressions during the search, only around were unique.
-This not only wastes computational resource but it also shows that, at some point, GP fails to explore different regions of the search space.
-Many authors observed improvements in the obtained solutions when applying any form of simplification during the search (Cao et al., 2023; Randall et al., 2022; Rivero et al., 2022; Seidyo Imai Aldeia et al., 2024) and, as a side effect, it stimulates the diversity of the population (Burlacu et al., 2019, 2020a).
+This not only wastes computational resources but it also shows that, at some point, GP fails to explore different regions of the search space.
+Many authors observed improvements in the obtained solutions when applying any form of simplification during the search (Cao et al., 2023; Randall et al., 2022; Rivero et al., 2022; Imai Aldeia et al., 2024) while also stimulating the diversity of the population (Burlacu et al., 2019, 2020a).
 Equality saturation has been used in the context of symbolic regression as a support tool to study the behavior of the search.
-Many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters (de Franca and Kronberger, 2023; Kronberger and de França, 2024).
+Many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters (de Franca and Kronberger, 2023; Kronberger and Olivetti de França, 2025).
 This redundancy can increase the chance of failing to correctly optimize such parameters, leading to sub-optimal solutions.
@@
 So far, the equality saturation technique was not used during the GP search to improve the quality of the solutions.
+Semantic similarity is frequently studied in the GP literature, either to improve the population diversity, the locality of the perturbation operators, or to understand the dynamics of GP search (Vanneschi et al., 2014).
+The semantic aware operators (Uy et al., 2010) introduce locality by replacing subtrees of an expression with semantically similar trees.
+Another approach is to combine two expressions by generating a random expression with a codomain in the range and creating the combined expression , ensuring a balance between the semantics of and (Moraglio et al., 2012).
+In (Ruberto et al., 2019), the authors introduce the equivalence function that determines whether two expressions are equivalent if the differences in their behavior in the semantic space is constant.
+Using this idea, they implement a filtering mechanism that rejects any offspring that is equivalent to any expression in the current population.
+Given the definition of mathematical equivalence, stating that , where is the variable domain, it is important to highlight that in the Semantic GP literature, semantic equivalence is often calculated using a limited number of data points (), which cannot guarantee equivalence, but is sufficient for an approximate measure of semantic similarity.
+Using equality saturation we can produce and store equivalent expressions without the need of evaluation, as explained in the next section.
+In this paper, we are concerned with semantic equality and the means to enforce the creation of unvisited expressions, regardless of locality.
 3. Equality saturation and e-graphs
@@
 Each solid box represents an e-node that contains a symbol of the expression.
-The dashed boxes group a set of e-nodes together and it is called an e-class, each e-class is assigned an e-class id (number in the bottom right of an e-class box).
-The main property of an e-class is that, no matter which e-node is chosen during the traversal, it will lead to an equivalent expression with all other e-nodes.
-We can see that in the middle box (e-class id ), if we follow through it will generate the expression and if we follow through it will generate .
+The dashed boxes, called e-class, group a set of e-nodes together.
+Each one of these e-classes is assigned an id (number in the bottom right of an e-class box).
+The main property of an e-class is that, no matter which e-node is chosen during the traversal, it will lead to an equivalent expression to all other e-nodes of the same e-class.
+Looking at the middle box (e-class id ), if we follow through it will generate the expression and if we follow through it will generate .
 The abstract description of the algorithm is very simple, though a concrete and optimal implementation requires the use of advanced techniques and data structures.
 The main idea is: i) match all the equivalence rules in the current state of the e-graph, ii) apply the rules creating new e-classes, iii) merge the equivalent e-classes, iv) repeat until saturation (i.e., no changes occur).
-The current state-of-the-art implementation (Willsey et al., 2021)222https://docs.rs/egg/latest/egg/, called egg333egg stands for e-graphs good, encodes the e-graph as a map of an e-node to an e-class, a map of the e-class ids to their structure, and a set of e-classes that still needs analysis to maintain consistency.
-The structure of an e-class contains the information about the e-nodes it contains, a list of the parent e-nodes, and additional information also referred to as semantic analysis.
+The data structure of an e-class stores the information about the e-nodes it contains, a list of the parent e-nodes, and additional information also referred to as semantic analysis.
 This implementation also maintains a database of patterns that allows the algorithm to efficiently match patterns inside the e-graph structure.
-Usually, the e-graph is used to represent a single program or expression and their equivalent forms.
-But, the structure can keep any number of expressions as long as we keep a list of the e-classes ids representing the root of each expression.
+The e-graph is commonly used to represent a single program or expression and their equivalent forms, when applying to simplification.
+But, the structure can keep any number of expressions as long as we keep a list of the e-classes ids that represents the root of each expression.
 For example, if we insert the expressions into the e-graph, we would end up with Fig.
 1a, minus the equivalent relations.
-As a result, we would keep the list representing the ids of the expressions we inserted.
+As a result, we would keep the list 222Underlined numbers represent e-class ids.
+representing the ids of the expressions we inserted.
 New expressions are added bottom-up.
-Starting from a terminal, the algorithm checks whether it already exists in the e-graph and returns its e-class id, if it does not exist, it creates a new id.
+Starting from a terminal, the algorithm checks whether it already exists in the e-graph returning its e-class id if it does, otherwise, it creates a new id.
 When adding an internal node, the algorithm first converts it to an e-node by replacing its children by their e-class ids and then it checks whether it already exists in the e-graph, returning the corresponding e-class id or a new one.
 In our example from Fig.
-1, if we try to add the expression , it would first retrieve the e-class ids for the terminals and , then it would return the e-class id corresponding to the e-node , where underlined numbers correspond to e-class ids.
+1, if we try to add the expression , it would first retrieve the e-class ids for the terminals and , then it would return the e-class id corresponding to the e-node .
 Finally, it would create a new e-class with id and the e-node .
 This mechanism allows us to compactly store a set of expressions and readily assert whether an expression already exists in the structure.
-In our implementation, we made a series of improvements to this data structure as we will describe in the next section.
 4. eggp: e-graph GP
 The proposed algorithm, eggp (e-graph genetic programming), follows the same structure as the traditional GP.
-The initial population is created using ramped half-and-half respecting a maximum size and maximum depth parameter (Koza, 1992) and, for a number of generations, it will choose two parents using tournament selection, apply the subtree crossover with probability followed by the subtree mutation with probability , when the offsprings replace the current population following a dominance criteria.
+The initial population of size is created using ramped half-and-half respecting a maximum size and maximum depth (Koza, 1992) and, for a number of generations, it will choose pair of parents using tournament selection, applying the subtree crossover with probability followed by the subtree mutation with probability , replacing the offsprings following a certain criteria.
 The key differences of eggp are:
-new solutions are inserted into the e-graph followed by one step of equality saturation to find and store some of the equivalent expressions of the new offspring.
-the current population is replaced by the set of individuals formed by: the Pareto front, the next front after excluding the first Pareto-front, and a selection of the last offspring at random until it reaches the desired population size.
+a single step of equality saturation is executed after inserting new expressions, merging equivalent expressions.
 the subtree crossover and mutation are modified to try to generate an unvisited expression.
-Apart from that, we do not make use of other mutation operators commonly used in the literature (Burlacu et al., 2020b; Cranmer, 2023) or advanced techniques to stimulate diversity such as island models.
-Notice that a single step of equality saturation will not guarantee the insertion of all equivalent expressions in the e-graph but, as we apply more iterations, the e-graph can grow exponentially large.
+Notice that a single step of equality saturation will not guarantee the insertion of all equivalent expressions in the e-graph but, if we apply more iterations, the e-graph can grow exponentially large.
 This issue is amplified by the fact that we are storing multiple expressions.
 As we will see in Sec.
-6, the single step seems to be sufficient to improve the results and the benefits of increasing the number of steps is a subject for a future research.
-Regarding the Pareto-front extracted from the entire history, this is equivalent to the traditional NSGA-II algorithm (Deb et al., 2000) as the dominance relation is transitive.
+6, the single step seems to be sufficient to improve the results and the benefits of increasing the number of steps is a subject for future research.
+We implemented single and multi-objective versions (called eggpso and eggpmo), with eggpso replacing the current population with the generated offspring and eggpmo replacing it by the set of individuals formed by: the Pareto front, the next front after excluding the first Pareto-front, and a selection of the last offspring at random until it reaches the desired population size.
 Keeping two ranks of dominance and filling up the remainder of the population with new expressions is meant to stimulate the combination of new expressions (exploration) while keeping the best fronts (exploitation).
-As a side-note, in this implementation we keep a database of generated expressions sorted by fitness and size (both objectives used in this work), so the Pareto-front can be retrieved in where is the number of extracted individuals.
+At the end of the execution, we can extract the Pareto-front from the entire history of the search, this is equivalent to the traditional NSGA-II algorithm (Deb et al., 2000) as the dominance relation is transitive.
+Moreover, we keep a database of generated expressions sorted by fitness and size (both objectives used in this work), so the Pareto-front can be retrieved in where is the number of extracted individuals.
 A new expression can be inserted into this structure in , where is the number of evaluated expressions so far.
-Finally, if we add the expression , equality saturation will generate the equivalent form (constants are replaced by parameters), this expression is stored in the database of expressions with size .
-4.1. Crossover and mutation
-Since the algorithm keeps the whole history of visited expression and its equivalent forms, we can exploit this information to increase the probability of generating unvisited expressions in the perturbation operators.
-For the crossover (Fig.
-2a), the main idea is that it first chooses a crossover point from the first parent and then it randomly chooses a subtree of the second parent which generates an unvisited expression upon combination.
-For the mutation operator (Fig.
-2b), it chooses a point of the tree to generate a new subtree and, after generating the new subtree, it checks whether it generates an unvisited expression.
-With a negative answer, it randomly changes the root node of this subtree from a choice of all the possible non-terminals that can ensure novelty.
-In the event of impossibility of generating a new expression, a random recombination is returned.
-To verify whether the combination of the original tree with the subtree replacing the -th node will generate an unvisited expression, we traverse the tree in pre-order traversal until we reach the node we want to replace (-th node).
-While traversing the tree, we build a list of -functions equivalent to a function that, given a tree , it creates a new tree with the operator of as the root, replacing one child with and keeping the other child of (if it is a binary operator).
-So, if the next step in the traversal is to the left, this function will build a tree rooted at with the left child being replaced by and the right child as , this is illustrated in Alg.
-1.
-Upon reaching the desired node, we traverse the tree upwards by checking whether the current node exists in the e-graph and, if the assertion is false, it will apply the first function of the list to , generating the parent e-node, and calling this same function recursively (Alg.
-2).
-Given that none of the ancestors exist in the e-graph, the algorithm confirms that this is an unvisited expression.
-The full implementation of eggp has nine hyperparameters: number of generations, population size, maximum expression size, likelihood function (MSE, Gaussian, Poisson, Bernoulli, ROXY (Lelli et al., 2017)), number of iterations and retries for the parameter optimization, probabilities of crossover and mutation, and the list of non-terminals.
-The algorithm is implemented in Haskell using the srtree444https://github.com/folivetti/srtree library for symbolic regression and with an adapted implementation of equality saturation based on hegg555https://github.com/alt-romes/hegg.
-The binaries are available at https://github.com/folivetti/srtree/releases/tag/v2.0.1.0.
+Finally, if we add the expression , equality saturation will generate the equivalent form (constants are replaced by parameters), storing it in the database of expressions with size .
+4.1. E-graph crossover and mutation
+The e-graph crossover and mutation operators exploits the information of the search history stored on the e-graph to increase the probability of generating an unvisited expression.
+The e-graph crossover (Fig.
+2a) involves two parents chosen with tournament selection and replaces a random subtree of the first parent with a random subtree sampled from a subset of all possible subtrees of the second parent.
+This subset is built such that, when replacing the chosen subtree of the first parent, it will generate an unvisited expression.
+In the event that this set is empty or the algorithm chooses not to perform the crossover (with probability ), it will return the unmodified first parent.
+The e-graph mutation (Fig.
+2b) is applied to the offspring of the crossover with a probability .
+Like the traditional subtree mutation, it replaces a random subtree of that solution with a randomly generated subtree using either the grow or full method (chosen at random).
+After the new expression is created, it is checked whether it already exists in the current e-graph.
+If it does, the node at the root of the generated subtree is exchanged by another node chosen at random from a subset of the symbols with the same arity.
+This subset is formed by all the symbols that would create an unvisited expression.
+When this subset is empty, the current mutated expression is returned.
+The full implementation of eggp has nine hyperparameters: number of generations, population size, maximum expression size, loss function (MSE, Gaussian, Poisson, Bernoulli, ROXY (Lelli et al., 2017)), number of iterations and retries for the parameter optimization, probabilities of crossover and mutation, and the list of non-terminals.
 5. Experiments
@@
 It also applies a simplification heuristic on a selection of the expressions.
-Both PySR and Operon uses multi-objective optimization optimizing accuracy and size as default.
-The research questions we want to address with the experiments are:
-What is the impact of increasing the probability of generating novelty, when compared to a minimalistic implementation such as tinyGP?
-How close does eggp get to the state-of-the-art without resorting to more advanced concepts such as specialized mutation operators, enforcing the placement of numerical parameters, and island model?
+Both PySR and Operon uses multi-objective optimization with accuracy and expression size as the default objectives.
 We should stress that we have kept eggp with only the subtree crossover and mutation to be directly comparable with tinyGP, thus measuring the benefits of this modification.
-We have fixed all the common hyperparameters to a standard value and performed a hyperparameter tuning with two different settings for each algorithm varying only whether to have a larger population with fewer iterations or a smaller population with more iterations.
-The only differences in settings are: for PySR we are using populations in its island model, so the size of each population is of the population size for the other algorithms, earlier experiments with the first set of benchmarks revealed that PySR performs significantly worse when using a single population; for Operon, we perform a maximum of optimization iterations instead of iterations with different starting points since it does not support multiple restarts for the parameter optimization; for eggp, of the training data is separated and used as a validation set to calculate the fitness, while the parameters are fitted using the remaining .
-Notice that tinyGP is the only single-objective approach and will be left out of the hypervolume comparison.
+We have fixed all the common hyperparameters to an empirically set of default values.
+The only differences in settings are: for PySR we are using populations in its island model, so the size of each population is of the population size for the other algorithms, earlier experiments revealed that PySR performs significantly worse when using a single population; for Operon, we perform a maximum of optimization iterations instead of iterations with different starting points since it does not support multiple restarts; for eggp, of the training data is separated and used as a validation set to calculate the fitness, while the parameters are fitted using the remaining .
 | Parameter | Value |
-| number of evaluations | |
+| pop. size / gens. / tourn. size | |
 | prob. mutation | |
@@
 | non-terminal set | |
-| pop. size / gens. / tourn. size | |
 | max depth | |
@@
 | optimization steps | ( for Operon) |
-For the first set of experiments, we picked four datasets from the original SRBench (de Franca et al., 2024) and split each one into training and test sets with a ratio of .
-We executed each configuration times and stored the test errors of each run.
-Once we have the optimal hyperparameters, we evaluate these algorithms in the benchmarks of the current version of SRBench666https://github.com/cavalab/srbench/discussions/174#discussioncomment-10285133 for the reduced SRBench track.
+For the first set of experiments, we evaluated each algorithm using the recently proposed333https://github.com/cavalab/srbench/discussions/174#discussioncomment-10285133 reduced SRBench benchmark.
 This set is supposed to be representative as it contains datasets with different characteristics.
@@
 Finally, we picked some real-world datasets from the literature corresponding to data from different fields, these data already have pre-determined training and test sets, as such we will run repetitions of each experiment.
-For every experiment, we store the final Pareto front and will report the performance plot, the AUC and average rank among all datasets and statistical test of the ranks, the average and standard deviation of the running time, the hypervolume of the Pareto front.
+For every experiment, we store the final Pareto front and will report the performance plot, the area under the curve (AUC), average rank among all datasets, statistical test of the ranks, the average and standard deviation of the running time.
 All algorithms were restricted to a single core to ensure equal conditions.
-Table 2 shows the datasets of each set of experiment with their corresponding number of data points, features and the chosen maximum size parameter.
-777The maximum size for Operon is set to of the maximum size because, internally, Operon will not count the scale coefficients of a terminal towards the model size.
-This factor enforces Operon to search on a similar search space as the other algorithms.
+Table 2 shows the datasets with their corresponding number of data points, features and the chosen maximum size parameter.
 | Name | Points | Features | max. size |
-| hyperparameter tuning | |||
-| 523_analcatdata_neavote | |||
-| 527_analcatdata_election2000 | |||
-| 635_fri_c0_250_10 | |||
-| 1029_LEV | |||
 | SRBench | |||
@@
 6. Results and Discussion
-6.1. Hyperparameter tuning
-Applying the Wilcoxon rank test to the obtained results using , we observed that only tinyGP expressed a significant difference in their results (-value ) favoring generations and population size .
-There was no significant difference observed in any algorithm, thus, we chose the configuration which found the best result for most datasets.
-With this criteria, PySR and Operon was set to generations and eggp with generations.
-6.2. New SRBench datasets
 In Fig.
-3 we can see the performance plots for each SRBench dataset of the best solution according to the highest on the training set.
+3 we can see the performance plots for the SRBench datasets considering the best solution of each run according to the highest on the training set.
 The x-axis of these plots represents the measured on the test set, and the y-axis shows the percentage of runs that the algorithm found an equal or larger than .
 The ideal algorithm would cover the whole area from to .
-As we can see from these plots, eggp always covers an area similar or better than the competing algorithms.
-Considering the area under the curve values (AUC), eggp covered the most area in out of the datasets, Operon in , PySR in and tinyGP in .
-It is also evident from these plots how eggp displays a significant improvement over tinyGP.
+From these plots we can see that in at least datasets (, and ) every algorithm reliably achieve the same .
+In other datasets (, , ) both versions of eggp maintain this reliability (i.e., achieves the same score in almost every execution) while the other algorithms either achieve a lower score or fails in some execution.
+The failing executions can be identified as , for dataset every algorithm fails between to of the times.
 In Table 3 we can see the ranks when considering the median of the test set and the AUC.
-We can see that, using this criteria, eggp is consistently ranked second, while Operon and PySR take turns in the first and third place.
+Using this criteria, eggpmo have an average rank of , while Operon comes next ranked on average.
+Considering eggpso and tinyGP, there is a slightly decrease in the average rank when using the proposed operators.
 The statistical test reveals that we can reject the null hypothesis when comparing to PySR with the alternative of having greater median rank.
-On the other hand, for the AUC values, we can see that eggp is greater than the other algorithms on average while rejecting the null hypotheses for each comparison.
-Unlike the median of the , the AUC is the average weighted by the probability of obtaining that value or greater, acting as a reliability measure of obtaining that value or greater.
-| dataset | eggp | Operon | PySR | tinyGP |
-| 192_vineyard | 1 | 4 | 3 | 2 |
-| 210_cloud | 1 | 4 | 2 | 3 |
-| 522_pm10 | 3 | 4 | 1 | 2 |
-| 557_analcatdata_apnea1 | 2 | 4 | 3 | 1 |
-| 579_fri_c0_250_5 | 2 | 1 | 3 | 4 |
-| 606_fri_c2_1000_10 | 1 | 2 | 3 | 4 |
-| 650_fri_c0_500_50 | 2 | 1 | 3 | 4 |
-| 678_visualizing_environmental | 1 | 4 | 2 | 3 |
-| 1028_SWD | 4 | 1 | 2 | 3 |
-| 1089_USCrime | 1 | 3 | 4 | 2 |
-| 1193_BNG_lowbwt | 2 | 1 | 3 | 4 |
-| 1199_BNG_echoMonths | 4 | 3 | 1 | 2 |
-| mean | 2.00 | 2.67 | 2.50 | 2.83 |
-| -value | ||||
-| 192_vineyard | 0.24 | 0.08 | 0.16 | 0.17 |
-| 210_cloud | 0.71 | 0.34 | 0.71 | 0.58 |
-| 522_pm10 | 0.16 | 0.14 | 0.19 | 0.18 |
-| 557_analcatdata_apnea1 | 0.75 | 0.40 | 0.58 | 0.68 |
-| 579_fri_c0_250_5 | 0.91 | 0.95 | 0.83 | 0.79 |
-| 606_fri_c2_1000_10 | 0.96 | 0.96 | 0.82 | 0.70 |
-| 650_fri_c0_500_50 | 0.87 | 0.92 | 0.83 | 0.40 |
-| 678_visualizing_environmental | 0.28 | 0.08 | 0.27 | 0.20 |
-| 1028_SWD | 0.38 | 0.39 | 0.39 | 0.36 |
-| 1089_USCrime | 0.68 | 0.59 | 0.59 | 0.59 |
-| 1193_BNG_lowbwt | 0.56 | 0.57 | 0.56 | 0.55 |
-| 1199_BNG_echoMonths | 0.40 | 0.30 | 0.43 | 0.35 |
-| mean | 0.58 | 0.48 | 0.53 | 0.46 |
-| -value |
-6.3. Real-world datasets
+On the other hand, for the AUC values, we can see that eggpmo is greater than the other algorithms on average while rejecting the null hypotheses for each comparison, except Operon.
+Unlike the median of the , the AUC is the average weighted by the probability of obtaining that value or greater, acting as a reliability measure.
+Also in this table, we can see that, on average, eggp (both versions) consistently return smaller models than the competing algorithms.
+There are two possible reasons for this behavior: i) as we apply equality saturation after inserting each expression into the e-graph, they can result in a simplified version of the inserted expression, ii) during the insertion, it automatically eliminates some of the redundant parameters, avoiding the issue reported in(de Franca and Kronberger, 2023).
+| dataset | eggpmo | eggpso | Operon | PySR | tinyGP |
+| 192 | |||||
+| 210 | |||||
+| 522 | |||||
+| 557 | |||||
+| 579 | |||||
+| 606 | |||||
+| 650 | |||||
+| 678 | |||||
+| 1028 | |||||
+| 1089 | |||||
+| 1193 | |||||
+| 1199 | |||||
+| mean | |||||
+| -value | 0.05 | 0.21 | 0.02 | 0.11 | |
+| 192 | |||||
+| 210 | |||||
+| 522 | |||||
+| 557 | |||||
+| 579 | |||||
+| 606 | |||||
+| 650 | |||||
+| 678 | |||||
+| 1028 | |||||
+| 1089 | |||||
+| 1193 | |||||
+| 1199 | |||||
+| mean | |||||
+| -value | 0.05 | 0.23 | 0.05 | 0.01 | |
+| 192 | |||||
+| 210 | |||||
+| 522 | |||||
+| 557 | |||||
+| 579 | |||||
+| 606 | |||||
+| 650 | |||||
+| 678 | |||||
+| 1028 | |||||
+| 1089 | |||||
+| 1193 | |||||
+| 1199 | |||||
+| mean |
 In Fig.
-4 we observe a similar behavior with eggp covering an area close to or better than the best competing algorithm.
-Considering AUC, it found out of best results, while PySR only , Operon and tinyGP .
-Again, considering the ranks on the median of the (Table 4), it consistently achieved second place, but in this set, Operon was ranked first for most of the datasets.
+4 we observe a similar behavior for the real-world datasets with eggp covering an area close to or better than the best competing algorithm.
+The largest difference was on the friction dataset in which eggpmo covered an area smaller th Operon.
+Considering AUC, eggpmo, Operon, eggpso obtained a similar average score and were ranked in this order.
+PySR and tinyGP obtained significatively worse average scores.
+Considering the ranks on the median of the (Table 4), the multi-objective version consistently achieved second place, but in this set, Operon was ranked first for most of the datasets.
 We should notice, though, that eggp results were always close to the best competing algorithm, while even Operon misbehaved in two datasets (flow and niku-2).
-When observing the statistical test results, we can conclude that there are no significant differences between the eggp and Operon, but the hypotheses of eggp being equivalent to PySR and tinyGP can be rejected.
-| dataset | eggp | Operon | PySR | tinyGP |
-| Chemical_1_tower | 2 | 1 | 4 | 3 |
-| Chemical_2_competition | 2 | 1 | 4 | 3 |
-| Flow_stress_phip0.1 | 1 | 4 | 2 | 3 |
-| Friction_dyn_one-hot | 2 | 1 | 4 | 3 |
-| Friction_stat_one-hot | 2 | 1 | 4 | 3 |
-| Nasa_battery_1_10min | 2 | 1 | 4 | 3 |
-| Nasa_battery_2_20min | 2 | 3 | 4 | 1 |
-| Nikuradse_1 | 2 | 1 | 4 | 3 |
-| Nikuradse_2 | 3 | 1 | 4 | 2 |
-| mean | 2.00 | 1.56 | 3.78 | 2.67 |
-| -value | ||||
-| Chemical_1_tower | 0.86 | 0.89 | 0.67 | 0.81 |
-| Chemical_2_competition | 0.58 | 0.66 | 0.07 | 0.37 |
-| Flow_stress_phip0.1 | 0.99 | 0.51 | 0.98 | 0.88 |
-| Friction_dyn_one-hot | 0.81 | 0.85 | 0.66 | 0.74 |
-| Friction_stat_one-hot | 0.77 | 0.82 | 0.30 | 0.65 |
-| Nasa_battery_1_10min | 0.96 | 0.96 | 0.85 | 0.93 |
-| Nasa_battery_2_20min | 0.97 | 0.95 | 0.97 | 0.98 |
-| Nikuradse_1 | 0.85 | 0.89 | 0.62 | 0.75 |
-| Nikuradse_2 | 0.98 | 0.83 | 0.95 | 0.95 |
-| mean | 0.87 | 0.82 | 0.67 | 0.79 |
-| -value |
-6.4. Hypervolume
-Table 5 shows the average and standard deviation of the hypervolume of the final Pareto fronts.
-For SRBench, eggp covered a similar area to the other algorithms.
-PySR returned the largest hypervolume in most of the SRBench datasets.
-For the real-world datasets, eggp covered the most hypervolume overall.
-| dataset | egraphGP | operon | PySR |
-| srbench | |||
-| 192 | |||
-| 210 | |||
-| 522 | |||
-| 557 | |||
-| 579 | |||
-| 606 | |||
-| 650 | |||
-| 678 | |||
-| 1028 | |||
-| 1089 | |||
-| 1193 | |||
-| 1199 | |||
-| real-world | |||
-| Chemical_1 | |||
-| Chemical_2 | |||
-| Flow_stress | |||
-| Friction_dyn | |||
-| Friction_stat | |||
-| Nasa_1 | |||
-| Nasa_2 | |||
-| Nikuradse_1 | |||
-| Nikuradse_2 | |||
-6.5. Computational time
-Regarding the runtime, since Operon is the fastest symbolic regression implementation, as noted in (Burlacu et al., 2020b), we calculated the ratio between the average runtime of each algorithm to Operon.
+When observing the statistical test results, we can conclude that there are no significant differences between the eggpmo and Operon, but the hypotheses of eggp being equivalent to PySR and tinyGP can be rejected.
+Regarding the model size, for this set of benchmarks in which we used a smaller maximum size, all algorithms returned a model size close to the maximum, on average.
+In this criteria, PySR consistently returned smaller models but with the expense of smaller accuracy as seen in the AUC results.
+| dataset | eggpmo | eggpso | Operon | PySR | tinyGP |
+| chemical 1 | |||||
+| chemical 2 | |||||
+| flow | |||||
+| friction | |||||
+| friction dyn | |||||
+| nasa 1 | |||||
+| nasa 2 | |||||
+| niku 1 | |||||
+| niku 2 | |||||
+| mean | |||||
+| -value | 0.02 | 0.93 | 0.00 | 0.01 | |
+| chemical 1 | |||||
+| chemical 2 | |||||
+| flow | |||||
+| friction | |||||
+| friction dyn | |||||
+| nasa 1 | |||||
+| nasa 2 | |||||
+| niku 1 | |||||
+| niku 2 | |||||
+| mean | |||||
+| -value | 0.00 | 0.85 | 0.00 | 0.00 | |
+| chemical 1 | |||||
+| chemical 2 | |||||
+| flow | |||||
+| friction | |||||
+| friction dyn | |||||
+| nasa 1 | |||||
+| nasa 2 | |||||
+| niku 1 | |||||
+| niku 2 | |||||
+| mean | |||||
+| mean |
+Regarding the computational runtime, since Operon is the fastest symbolic regression implementation, as noted in (Burlacu et al., 2020b), we calculated the ratio between the average runtime of each algorithm to Operon.
 Fig.
@@
 We should stress that all algorithms were constrained to run with a single thread, thus both Operon and PySR runtime could be smaller when exploiting multi-threading.
+6.1. Additional considerations
+The main limitations of these experiments lie in the use of default or reasonable values for the hyperparameters.
+We should notice that eggp contains hyperparameters that should be fine-tuned to obtain optimal results in a practical scenario.
+Meanwhile, PySR contains about hyperparameters that may affect the algorithm performance444we are not considering hyperparameters that can be set using prior knowledge.
+and Operon contains about hyperparameters.
+With careful experimentation, both PySR and Operon could have achieved similar results to those obtained with our approach.
+Having said that, eggp is a step forward to a parameterless experience in SR implementations, where the user only needs to set hyper-parameters that are intuitive w.r.t.
+their behavior.
+For example, increasing the number of evaluations will never make the results worse, unlike crossover and mutation rates which behaves unpredictably.
+Another feature of eggp is the ability to export the current state of the e-graph into a file and load it when starting a new search.
+When loading a previous e-graph, eggp will resume the search by first extracting an initial population from that e-graph.
+In such case, eggpmo will recover the current Pareto front and eggpso will sample random solutions.
+The search can be resumed with different hyper-parameters, so the user can include new non-terminals or increase the maximum size.
+The e-graph file can also be used with the exploration tool rEGGression(de Franca and Kronberger, 2025) that allows the user to explore the history of solution and retrieve a regression model outside of the Pareto front.
+6.2. Data availability
+The algorithm is implemented in Haskell using the srtree555https://github.com/folivetti/srtree library for symbolic regression and equality saturation.
+The binaries and source code of eggp and tinyGP used in this paper are available at https://github.com/folivetti/srtree/releases/tag/v2.0.1.0 and all the datasets, scripts, and results to replicate the experiments are available at https://github.com/folivetti/eggp_paper_GECCO.
 7. Conclusions
-In this paper we explored the use of e-graphs and equality saturation as a mechanism to keep track of the history of the search engine of symbolic regression and exploiting its pattern matching capabilities to propose a variation to the traditional subtree crossover and mutation that increases the probability of generating novel expressions.
-The e-graph data structure efficiently stores and queries for parts of expressions.
-Exploiting this capability, we modified the subtree operators to only sample subtrees that would form unvisited expressions.
+In this paper we explored the use of e-graphs and equality saturation as a mechanism to keep track of the history of the search engine for symbolic regression and to exploit its pattern matching capabilities to propose a variation to the traditional subtree crossover and mutation that increases the probability of generating novel expressions.
+The e-graph data structure compactly stores shared elements of a set of expressions and their equivalence relationships, and efficiently queries for parts of expressions.
+Exploiting this capability, we modified the subtree operators to only sample subtrees that would generate unvisited expressions.
 The expectation is that this simple modification would render a significant improvement in the search procedure.
-We tested the proposed algorithm, called eggp, in different benchmarks from the literature and compared with a simple GP with the original subtree operators, and two state-of-the-art algorithms, PySR and Operon.
+We tested the proposed algorithm, called eggp, in different benchmarks from the literature and compared with a simple GP using the original subtree operators, and two state-of-the-art algorithms, PySR and Operon.
 The results showed that the modified operators are capable of improving the performance of a simple GP to compete with the state-of-the-art.
 The main highlight of this approach is that it consistently performs equal or better than the best competing approach.
-The main limitations of these experiments lie in the use of default or reasonable values for the hyperparameters.
-We should notice that eggp contains eight parameters that should be adjusted accordingly in a practical scenario.
-Meanwhile, PySR contains about hyperparameters that may affect the algorithm performance888we are not considering hyperparameters that expect prior knowledge.
-and Operon contains about hyperparameters.
-With careful experimentation, both PySR and Operon could possibly achieve similar results.
 Regarding the runtime, eggp is consistently faster than PySR but significantly slower than Operon.
@@
 This allowed us to obtain a more robust algorithm delivering better performance more reliably than state-of-the-art implementations.
-As for the next steps, this same modifications can be applied to any other mutation operator used by the state-of-the-art algorithms.
+As for the next steps, this same modifications can be applied to any other operator used by the state-of-the-art algorithms.
 Not only that, but the e-graph opens up many new possibilities for improving the search as it allows us to query expressions with a combination of properties, which can translate to diversity-preserving population and easy to integrate prior knowledge (Kronberger et al., 2022; Reuter et al., 2024).
 In addition, the storage of the search history allows us to analyze the learned building blocks and exploit this information to generate new solutions.
-999The authors proudly made no use of LLMs for this work.
 Acknowledgements.
@@
 is supported by the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology, the Federal Ministry for Labour and Economy, and the regional government of Upper Austria within the COMET project ProMetHeus (904919) supported by the Austrian Research Promotion Agency (FFG).
-No LLMs were used for this work.References
+The authors proudly made no use of LLMs for this work.References
 - (1)
@@
 Springer, 65–86.
-- Bartlett et al. (2023) Deaglan J Bartlett, Harry Desmond, and Pedro G Ferreira. 2023. Exhaustive symbolic regression. IEEE Transactions on Evolutionary Computation (2023).
+- Bartlett et al.
+(2024) Deaglan J.
+Bartlett, Harry Desmond, and Pedro G.
+Ferreira.
+2024.
+Exhaustive Symbolic Regression.
+IEEE Transactions on Evolutionary Computation 28, 4 (2024), 950–964.
+https://doi.org/10.1109/TEVC.2023.3280250 arXiv:2211.11461 [astro-ph.CO]
 - Burlacu et al.
@@
 In 2019 IEEE Congress on Evolutionary Computation (CEC).
-IEEE, 2175–2182.
+2175–2182.
+https://doi.org/10.1109/CEC.2019.8790162
 - Burlacu et al.
@@
 Hash-Based Tree Similarity and Simplification in Genetic Programming for Symbolic Regression.
-In Computer Aided Systems Theory–EUROCAST 2019: 17th International Conference, Las Palmas de Gran Canaria, Spain, February 17–22, 2019, Revised Selected Papers, Part I 17.
-Springer, 361–369.
+In Computer Aided Systems Theory – EUROCAST 2019, Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia (Eds.).
+Springer International Publishing, Cham, 361–369.
 - Burlacu et al.
@@
 2020b.
-Operon C++ an efficient genetic programming framework for symbolic regression.
-In Proceedings of the 2020 Genetic and Evolutionary Computation Conference Companion.
-1562–1570.
+Operon C++: an efficient genetic programming framework for symbolic regression.
+In Proceedings of the 2020 Genetic and Evolutionary Computation Conference Companion (Cancún, Mexico) (GECCO ’20).
+Association for Computing Machinery, New York, NY, USA, 1562–1570.
+https://doi.org/10.1145/3377929.3398099
 - Cao et al.
-([n.
-d.]) Haotian Cao, Garrett W Merz, Kyle Cranmer, and Gary Shiu.
-[n.
-d.].
+(2024) Haotian Cao, Garrett W Merz, Kyle Cranmer, and Gary Shiu.
+2024.
 Learning Conformal Field Theory with Symbolic Regression: Recovering the Symbolic Expressions for the Energy Spectrum.
-([n.
-d.]).
+In NeurIPS 2024 Workshop: Machine Learning and the Physical Sciences.
 - Cao et al.
@@
 1064–1072.
+- de Franca and Kronberger (2025) Fabricio Olivetti de Franca and Gabriel Kronberger.
+2025.
+rEGGression: an Interactive and Agnostic Tool for the Exploration of Symbolic Regression Models.
+In Proceedings of the Genetic and Evolutionary Computation Conference (Malaga, Spain) (GECCO ’25).
+Association for Computing Machinery, New York, NY, USA.
+https://doi.org/10.1145/3712256.3726385 arXiv:2501.17859 [cs.LG]
 - de Franca et al.
-(2024) Fabricio O de Franca, Marco Virgolin, M Kommenda, MS Majumder, M Cranmer, G Espada, L Ingelse, A Fonseca, M Landajuela, B Petersen, et al.
+(2024) F.
+O.
+de Franca, M.
+Virgolin, M.
+Kommenda, M.
+S.
+Majumder, M.
+Cranmer, G.
+Espada, L.
+Ingelse, A.
+Fonseca, M.
+Landajuela, B.
+Petersen, R.
+Glatt, N.
+Mundhenk, C.
+S.
+Lee, J.
+D.
+Hochhalter, D.
+L.
+Randall, P.
+Kamienny, H.
+Zhang, G.
+Dick, A.
+Simon, B.
+Burlacu, Jaan Kasak, Meera Machado, Casper Wilstrup, and W.
+G.
+La Cavaz.
 2024.
 SRBench++: Principled Benchmarking of Symbolic Regression With Domain-Expert Interpretation.
-IEEE transactions on evolutionary computation (2024).
+IEEE Transactions on Evolutionary Computation (2024), 1–1.
+https://doi.org/10.1109/TEVC.2024.3423681
 - Deb et al.
-(2000) Kalyanmoy Deb, Samir Agrawal, Amrit Pratap, and Tanaka Meyarivan.
+(2000) Kalyanmoy Deb, Samir Agrawal, Amrit Pratap, and T.
+Meyarivan.
 2000.
 A Fast Elitist Non-dominated Sorting Genetic Algorithm for Multi-objective Optimization: NSGA-II.
-In Parallel Problem Solving from Nature PPSN VI: 6th International Conference Paris, France, September 18–20, 2000 Proceedings 6.
-Springer, 849–858.
-- Ebner et al. (2001) Marc Ebner, Mark Shackleton, and Rob Shipman. 2001. How neutral networks influence evolvability. Complex. 7, 2 (Nov. 2001), 19–33. https://doi.org/10.1002/cplx.10021
+In Parallel Problem Solving from Nature PPSN VI, Marc Schoenauer, Kalyanmoy Deb, Günther Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel (Eds.).
+Springer Berlin Heidelberg, Berlin, Heidelberg, 849–858.
+- Ebner et al.
+(2001) Marc Ebner, Mark Shackleton, and Rob Shipman.
+2001.
+How neutral networks influence evolvability.
+Complexity 7, 2 (2001), 19–33.
+https://doi.org/10.1002/cplx.10021 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1002/cplx.10021
 - Gustafson et al.
-(2005) Steven Gustafson, Edmund K Burke, and Natalio Krasnogor.
+(2005) S.
+Gustafson, E.K.
+Burke, and N.
+Krasnogor.
 2005.
@@
 1.
-IEEE, 912–919.
-- Hu and Banzhaf (2018) Ting Hu and Wolfgang Banzhaf. 2018. Neutrality, robustness, and evolvability in genetic programming. Genetic Programming Theory and Practice XIV (2018), 101–117.
+912–919 Vol.1.
+https://doi.org/10.1109/CEC.2005.1554780
+- Hu and Banzhaf (2018) Ting Hu and Wolfgang Banzhaf.
+2018.
+Neutrality, Robustness, and Evolvability in Genetic Programming.
+Springer International Publishing, Cham, 101–117.
+https://doi.org/10.1007/978-3-319-97088-2_7
 - Hu et al.
@@
 Phenotype Search Trajectory Networks for Linear Genetic Programming.
-In European Conference on Genetic Programming (Part of EvoStar).
-Springer, 52–67.
-- Keller and Banzhaf (1999) Robert E Keller and Wolfgang Banzhaf.
+In Genetic Programming, Gisele Pappa, Mario Giacobini, and Zdenek Vasicek (Eds.).
+Springer Nature Switzerland, Cham, 52–67.
+- Imai Aldeia et al.
+(2024) Guilherme Seidyo Imai Aldeia, Fabrício Olivetti De França, and William G.
+La Cava.
+2024.
+Inexact Simplification of Symbolic Regression Expressions with Locality-sensitive Hashing.
+In Proceedings of the Genetic and Evolutionary Computation Conference (Melbourne, VIC, Australia) (GECCO ’24).
+Association for Computing Machinery, New York, NY, USA, 896–904.
+https://doi.org/10.1145/3638529.3654147
+- Keller and Banzhaf (1999) Robert E.
+Keller and Wolfgang Banzhaf.
 1999.
 The evolution of genetic code in Genetic Programming.
-In Proceedings of the Genetic and Evolutionary Computation Conference, Vol.
-2.
-Morgan Kaufmann Orlando, 1077–1082.
+In Proceedings of the 1st Annual Conference on Genetic and Evolutionary Computation - Volume 2 (Orlando, Florida) (GECCO’99).
+Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 1077–1082.
 - Koza (1992) John R. Koza. 1992. Genetic Programming: On the Programming of Computers by Means of Natural Selection. MIT Press, Cambridge, MA, USA.
 - Kronberger et al. (2024a) Gabriel Kronberger, Bogdan Burlacu, Michael Kommenda, Stephan M. Winkler, and Michael Affenzeller. 2024a. Symbolic Regression. Chapman & Hall / CRC Press.
-- Kronberger and de França (2024) Gabriel Kronberger and Fabrício Olivetti de França.
-2024.
-Effects of reducing redundant parameters in parameter optimization for symbolic regression using genetic programming.
-Journal of Symbolic Computation (2024), 102413.
 - Kronberger et al.
-(2022) Gabriel Kronberger, Fabricio Olivetti de França, Bogdan Burlacu, Christian Haider, and Michael Kommenda.
+(2022) G.
+Kronberger, F.
+O.
+de Franca, B.
+Burlacu, C.
+Haider, and M.
+Kommenda.
 2022.
 Shape-Constrained Symbolic Regression—Improving Extrapolation with Prior Knowledge.
-Evolutionary Computation 30, 1 (2022), 75–98.
+Evolutionary Computation 30, 1 (03 2022), 75–98.
+https://doi.org/10.1162/evco_a_00294
 - Kronberger et al.
@@
 Springer Nature Switzerland, Cham, 273–289.
+- Kronberger and Olivetti de França (2025) Gabriel Kronberger and Fabrício Olivetti de França.
+2025.
+Effects of reducing redundant parameters in parameter optimization for symbolic regression using genetic programming.
+Journal of Symbolic Computation 129 (2025), 102413.
+https://doi.org/10.1016/j.jsc.2024.102413
 - La Cava et al.
@@
 Advances in neural information processing systems 2021, DB1 (2021), 1.
-- Lamb and Harrington ([n. d.]) Kara D Lamb and Jerry Y Harrington. [n. d.]. Discovering How Ice Crystals Grow Using Neural ODE’s and Symbolic Regression. ([n. d.]).
 - Lelli et al.
@@
 https://doi.org/10.1109/TEVC.2006.871253
+- Moraglio et al.
+(2012) Alberto Moraglio, Krzysztof Krawiec, and Colin G Johnson.
+2012.
+Geometric semantic genetic programming.
+In International Conference on Parallel Problem Solving from Nature.
+Springer, 21–31.
 - Murty and Kabadi (1985) Katta G Murty and Santosh N Kabadi. 1985. Some NP-complete problems in quadratic and nonlinear programming. Technical Report.
@@
 Expert Systems with Applications 198 (2022), 116712.
+- Ruberto et al. (2019) Stefano Ruberto, Leonardo Vanneschi, and Mauro Castelli. 2019. Genetic programming with semantic equivalence classes. Swarm and evolutionary computation 44 (2019), 453–469.
 - Russeil et al.
@@
 961–970.
-- Seidyo Imai Aldeia et al.
-(2024) Guilherme Seidyo Imai Aldeia, Fabricio Olivetti de Franca, and William G La Cava.
-2024.
-Inexact Simplification of Symbolic Regression Expressions with Locality-sensitive Hashing.
-arXiv e-prints (2024), arXiv–2404.
+- Schmidt and Lipson (2009) Michael Schmidt and Hod Lipson. 2009. Distilling free-form natural laws from experimental data. science 324, 5923 (2009), 81–85.
 - Sipper (2019) M. Sipper. 2019. Tiny Genetic Programming in Python. https://github.com/moshesipper/tiny_gp.
@@
 - Udrescu and Tegmark (2020) Silviu-Marian Udrescu and Max Tegmark. 2020. AI Feynman: A physics-inspired method for symbolic regression. Science Advances 6, 16 (2020), eaay2631.
-- Virgolin and Pissis ([n. d.]) Marco Virgolin and Solon P Pissis. [n. d.]. Symbolic Regression is NP-hard. Transactions on Machine Learning Research ([n. d.]).
+- Uy et al.
+(2010) Nguyen Quang Uy, Michael O’Neill, Nguyen Xuan Hoai, Bob Mckay, and Edgar Galván-López.
+2010.
+Semantic similarity based crossover in GP: The case for real-valued function regression.
+In Artifical Evolution: 9th International Conference, Evolution Artificielle, EA, 2009, Strasbourg, France, October 26-28, 2009.
+Revised Selected Papers 9.
+Springer, 170–181.
+- Vanneschi et al. (2014) Leonardo Vanneschi, Mauro Castelli, and Sara Silva. 2014. A survey of semantic methods in genetic programming. Genetic Programming and Evolvable Machines 15 (2014), 195–214.
+- Virgolin and Pissis (2022) Marco Virgolin and Solon P Pissis. 2022. Symbolic Regression is NP-hard. Transactions on Machine Learning Research (2022). https://openreview.net/forum?id=LTiaPxqe2e
 - Willsey et al.